# Nonlinear Schrödinger-Poisson systems in dimension two: the zero mass case

### v1

@@ 2.1. The space $\boldsymbol{D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})}$
$\|\tau_{h}u-u\|^{q}_{q}$ $=\int_{\mathbb{R}^{2}}|u(x+h)-u(x)|^{q}\,{\rm d}x\leq\left(\int_{\mathbb{R}^{2}}|u(x+h)-u(x)|^{2(q-1)}\,{\rm d}x\right)^{\frac{1}{2}}\|\tau_{h}u-u\|_{2}$
@@
2.2. Useful theorems and inequalities
Since our nonlinearities are of exponential growth, we recall here some important functional inequalities that we will exploit to control them. The first result is a generalised Cao’s inequality in $D^{1,2}(\mathbb{R}^{2})\cap L^{q}(\mathbb{R}^{2})$ taken from [28, Theorem 1.1], which is a particular case of the very general result obtained therein. Then, we report a Pohožaev-Trudinger-Moser inequality with logarithmic weight in $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ from [^1] Note that the space $H^{1}_{\omega}(\mathbb{R}^{2})$ in [14] coincides with our $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$. [14].
@@ 2.2. Useful theorems and inequalities
Let $q\geq p$ and $\lambda<4\pi$. Then there exists a constant $C:=C(p,q,\lambda)>0$ such that for all $u\in D^{1,2}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2})$ with $\|\nabla u\|_{2}\leq 1$ there holds
(2.4) $\int_{\mathbb{R}^{2}}{\rm e}^{\lambda u^{2}}|u|^{q}\,{\rm d}x\leq C\|u\|^{p}_{p}\,.$
If $\lambda\geq 4\pi$, (2.4) remains true but the constant $C$ is not uniform in $u$.
@@
Let $g$ satisfy assumptions $(f_{1})$ and either $(f_{sc})$ or $(f_{c})$, and $G(s)=\int_{0}^{s}g(t)\,{\rm d}t$. Then, the space $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ embeds into the weighted Orlicz space $L_{G}(\mathbb{R}^{2},\omega\,{\rm d}x)$, namely
$\int_{\mathbb{R}^{2}}G(\alpha|u|)\ln(b+|x|)\,{\rm d}x<+\infty$
for any $u\in D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ and any $\alpha>0$. Furthermore, for any $\alpha\leq 1/\sqrt{p}$ one has
Corollary 2.4 ([14], Corollary 3.4).
@@
Throughout the paper, we will make great use of the following well-known results: the Hardy-Littlewood-Sobolev inequality (see [31, Theorem 4.3]), and the Gagliardo-Nirenberg inequality (see [37, Formula (2.2)]).
@@
Let $N,m\in\mathbb{N}$ and $u\in L^{p}(\mathbb{R}^{N})$ such that $D^{m}u\in L^{r}(\mathbb{R}^{N})$, where $q,r\in[1,+\infty]$. Then, there exists a constant $C:=C(N,m,j,p,r,\theta)>0$ such that
(2.5) $\|D^{j}u\|_{q}\leq C\|D^{m}u\|_{r}\|u\|_{p},$
where $j$ satisfies
$\frac{1}{q}=\frac{j}{N}+\theta\left(\frac{1}{r}-\frac{m}{N}\right)+\frac{1-\theta}{p},\quad\mbox{ for every }j\mbox{ such that }\frac{j}{m}\leq\theta\leq 1.$
In particular:
  • if $j=0$, $rm<N$ and $p=+\infty$, then we need that either $u$ vanishes at infinity or $u\in L^{s}(\mathbb{R}^{N})$ for some $s\in(0,+\infty)$;
  • if $r\in(1,+\infty)$ and $m-j-N/r$ is a nonnegative integer, then (2.5) holds only for $\theta$ satisfying $\frac{j}{m}\leq\theta<1$.
We end this section by recalling an abstract result from [25, Proposition 3.1] (see also [48, Theorem 2.8] for the version with the Palais-Smale condition), which will be needed to prove the existence of bounded Cerami sequences.
@@
then, for every $\varepsilon\in\left(0,\frac{c-a}{a}\right)$, $\delta>0$, and $\gamma\in\widetilde{\Gamma}$, with $\sup_{u\in M}\Psi(\gamma(u))\leq c+\varepsilon$, there exists $u\in\widetilde{X}$ such that
@@ 2.3. Consequences of the assumptions
To end this Section, let us point out some immediate consequences of ($f_{0}$)-($f_{5}$) which will be of use in our analysis, together with some comments in this regard:
@@
    while, if $(f_{sc})$ is replaced by $(f_{c})$, the upper bound changes as follows: for any $r>0$, $\alpha>\alpha_{0}$ and $s_{0}>1$ there is $C>0$ such that such that
    (2.7) $0\leq F(s)\leq C\cdot\left\{\begin{array}[]{ll}s^{p}&\mbox{for }s\leq s_{0}\,,\\ s^{r}{\rm e}^{\alpha s^{2}}&\mbox{for }s>s_{0}\,,\end{array}\right.\quad\hbox{and }\quad 0\leq f(s)\leq C\cdot\left\{\begin{array}[]{ll}s^{p-1}&\mbox{for }s\leq s_{0}\,,\\ s^{r}{\rm e}^{\alpha s^{2}}&\mbox{for }s>s_{0}\,;\end{array}\right.$
@@
    ($f_{5}$) is related to the well-known de Figueiredo-Miyagaki-Ruf condition [23] and is crucial in order to estimate the mountain pass level and gain compactness, see Lemma 4.5. We note here that such an assumption, which goes back to [24], avoids the prescription of a global lower-bound on $F$ of the kind (2.8) but with $C$ large enough: it is widely used in the literature but is not of practical verification. A condition similar to ($f_{5}$) appears also e.g. in [14, 7, 17, 40].
@@
    Examples of nonlinearities which satisfy the assumptions of Theorem 1.2 are $F(s)=s^{p}$ and $F(s)=s^{p}{\rm e}^{s}$ with $p>2$ (concerning ($f_{sc}$)), and $F(s)=s^{p}\chi_{\{s<1\}}(s)+s^{q}{\rm e}^{s^{2}}\chi_{\{s\geq 1\}}(s)$ with $q>-2$ (concerning ($f_{c}$)).

### v2

@@ 2.1. The space $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$
$\|\tau_{h}u-u\|^{q}_{q}$ $=\int_{\mathbb{R}^{2}}|u(x+h)-u(x)|^{q}\,{\rm d}x$
$\leq\left(\int_{\mathbb{R}^{2}}|u(x+h)-u(x)|^{2(q-1)}\,{\rm d}x\right)^{\frac{1}{2}}\|\tau_{h}u-u\|_{2}$
@@
2.2. Functional inequalities in log-weighted Lebesgue spaces
Since our nonlinearities exhibit exponential growth, we present here some important functional inequalities, including newly established ones, to effectively handle them. The first result is a generalised Cao’s inequality in $D^{1,2}(\mathbb{R}^{2})\cap L^{q}(\mathbb{R}^{2})$ taken from [28], which is a particular case of the very general result obtained therein.
@@ 2.2. Functional inequalities in log-weighted Lebesgue spaces
Let $q\geq p$ and $\alpha<4\pi$. Then there exists a constant $C:=C(p,q,\alpha)>0$ such that for all $u\in D^{1,2}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2})$ with $\|\nabla u\|_{2}\leq 1$ there holds
(2.3) $\int_{\mathbb{R}^{2}}|u|^{q}\,{\rm e}^{\alpha u^{2}}\,{\rm d}x\leq C\|u\|^{p}_{p}\,.$
If $\alpha\geq 4\pi$, (2.3) remains true but the constant $C$ is not uniform in $u$.
Next, we recall a Pohožaev-Trudinger inequality with logarithmic weight in $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ from [14].
@@
Let $g$ satisfy assumptions $(f_{1})$ and either $(f_{sc})$ or $(f_{c})$ with exponent $\alpha_{0}$, and $G(s)=\int_{0}^{s}g(t)\,{\rm d}t$. Then, the space $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ embeds into the weighted Orlicz space $L_{G}(\mathbb{R}^{2},\omega\,{\rm d}x)$, namely
(2.4) $\int_{\mathbb{R}^{2}}G(\alpha|u|)\ln(b+|x|)\,{\rm d}x<+\infty$
for any $u\in D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ and any $\alpha>0$.
Reasoning as in [14, Corollary 3.4], from Theorem 2.3 it is easy to obtain the following:
Corollary 2.4.
@@
For our purposes, a uniform estimate in the spirit of Moser for the inequality (2.4) will be essential. In [14, Theorem 3.3] the authors obtained it under the condition $\alpha\leq\frac{2\pi}{\alpha_{0}\sqrt{p}}$ by relying on Ruf’s inequality. Here, we need to improve this threshold, in particular avoiding the dependence on $p$. In light of assumption ($f_{c}$), this is accomplished once the following Moser-type result is achieved, which extends to the space $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ the corresponding result in [47] for the case $p=2$.
Theorem 2.5.
For any $\alpha<4\pi$
$\sup_{\|u\|^{2}\leq 1}\int_{\mathbb{R}^{2}}|u|^{p}{\rm e}^{\alpha u^{2}}\ln(b+|x|)\,{\rm d}x<+\infty\,.$
The inequality is sharp, in the sense that, for any $\alpha>4\pi$
$\sup_{\|u\|^{2}\leq 1}\int_{\mathbb{R}^{2}}|u|^{p}{\rm e}^{\alpha u^{2}}\ln(b+|x|)\,{\rm d}x=+\infty\,.$
Proof.
The proof follows by suitably adapting the arguments in [47]. The main tool is a a transformation which relates functions in the weighted space $D^{1,2}L^{p}_{w}(\mathbb{R}^{2})$ to functions in the unweighted space $D^{1,2}_{0}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2})$, based upon a change of variables acting only on the radial part of $x$. The price to pay is a dilation term in the Dirichlet norm, whose effect is lower and lower as $|x|\to+\infty$ due to the logarithmic growth of the weight. This property is the key tool that allows to retain the sharp threshold $4\pi$, and it is the only one which needs a proper delicate modification when passing from [47] to our case. Let us denote by $T:\mathbb{R}^{2}\to\mathbb{R}^{2}$ the change of variable
$s=T(r)=\sqrt{2\int_{0}^{r}\rho\ln(b+\rho)\,{\rm d}\rho}\,.$
Setting
$v(y):=u(x),\ \ \ \ \ \hbox{namely, }\ \ v(y)=u\left(T^{-1}(|y|)\cos\theta,T^{-1}(|y|)\sin\theta\right),$
one can easily verify that that the map
$\mathcal{T}:D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$$\to$ $D^{1,2}_{0}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2})$
$u$ $\mapsto$ $v$
is an invertible, continuous map, with continuous inverse map too. Let $\chi(|x|)$ be a smooth, radial cut-off function
$\chi(|x|)=\begin{cases}1&\hbox{if }|x|<1\,,\\ 0&\hbox{if }|x|>2\,,\end{cases}$
and $\xi(|x|):=1-\chi(|x|)$; then scale $\xi$ as follows
$\xi_{\eta}(|x|):=\xi(\eta|x|)\quad\mbox{for}\ \eta\in(0,1)\,,$
and define $u_{\eta}:=u\cdot\xi_{\eta}$ for any $u\in X$ with $\|u\|\leq 1$; $u_{\eta}$ is then supported in $B_{1/\eta}^{c}$. We have
(2.5) $\int_{\mathbb{R}^{2}}|\nabla u_{\eta}|^{2}$ $\leq\|\nabla u\|_{2}^{2}+\left(\int_{u^{2}\leq\eta^{\frac{2}{p-2}}}|\nabla\xi_{\eta}|^{2}u^{2}+\int_{u^{2}>\eta^{\frac{2}{p-2}}}|\nabla\xi_{\eta}|^{2}u^{2}\right)$
$\quad+2\left(\int_{u^{2}\leq\eta^{\frac{2}{p-2}}}u\,\xi_{\eta}\nabla u\nabla\xi_{\eta}+\int_{u^{2}>\eta^{\frac{2}{p-2}}}u\,\xi_{\eta}\nabla u\nabla\xi_{\eta}\right)$
$\leq\|\nabla u\|_{2}^{2}+\eta^{\frac{2}{p-2}}\|\nabla\chi\|_{2}^{2}+\eta^{-1}\int_{u^{2}>\eta^{\frac{2}{p-2}}}|\nabla\xi_{\eta}|^{2}|u|^{p}$
$\quad+2\left(\eta^{\frac{1}{p-2}}\|\nabla u\|_{2}\|\xi_{\eta}\nabla\xi_{\eta}\|_{2}+\left[\int_{u^{2}>\eta^{\frac{2}{p-2}}}|u|^{2}\right]^{\frac{1}{2}}\|\nabla u\|_{2}\|\nabla\xi_{\eta}\|_{\infty}\right)$
$\leq\|\nabla u\|_{2}^{2}+\eta^{\frac{2}{p-2}}\|\nabla\chi\|_{2}^{2}+\eta\|\nabla\chi\|_{\infty}^{2}\int_{u^{2}>\eta^{\frac{2}{p-2}}}|u|^{p}$
$\quad+2\left(\eta^{\frac{1}{p-2}}\|\nabla u\|_{2}\|\nabla\chi\|_{2}+\sqrt{\eta}\|u\|_{\ast,p}^{\frac{p}{2}}\|\nabla u\|_{2}\|\nabla\chi\|_{\infty}\right)$
so that, if $\tau:=\min\big{\{}\frac{1}{2},\frac{1}{p-2}\big{\}}$,
(2.6) $\|u_{\eta}\|^{2}\leq 1+C\eta^{\tau}$
for some positive fixed constant $C$ independent of $\eta$ and for any $u\in X$ such that $\|u\|\leq 1$. The proof now follows as in [47]. Let $v_{\eta}$ be defined as
$\ u_{\eta}(r\cos\theta,r\sin\theta)=v_{\eta}\left(T(r)\cos\theta,T(r)\sin\theta\right);$
observing that
$\|\nabla v_{\eta}\|_{2}^{2}+\|v_{\eta}\|_{p}^{\frac{2}{p}}<\left(1+\frac{1}{2\ln(1+\frac{1}{\eta})}\right)\|u_{\eta}\|^{2}\leq\left(1+\frac{1}{2\ln(1+\frac{1}{\eta})}\right)(1+C\eta^{\tau})$
by (2.6) we can fix $\eta$ such that
$\left(1+\frac{1}{2\ln(1+\frac{1}{\eta})}\right)(1+C\eta^{\tau})<\frac{4\pi}{\alpha}\,,$
so that $\alpha\|v_{\eta}\|^{2}<4\pi$, and apply Theorem 2.2 to conclude.
The sharpness can be proved considering the same sequence of radial functions introduced in [47]
$u_{n}(x)=\frac{1}{\sqrt{2\pi}}\left\{\begin{array}[]{ll}\displaystyle\sqrt{\delta_{n}\ln n}&\displaystyle 0\leq|x|\leq\frac{R_{n}}{n}\,,\\ \\ \displaystyle\frac{\sqrt{\delta}_{n}}{\sqrt{\ln n}}\ln\left(\frac{R_{n}}{|x|}\right)&\displaystyle\frac{R_{n}}{n}<|x|\leq R_{n}\,,\end{array}\right.$
where
$R_{n}:=\frac{\sqrt{\ln n}}{\ln\ln n}\to+\infty\,,\quad\delta_{n}:=1-\frac{\ln\ln n}{4\ln n}\to 1^{-},\quad\text{ as }n\to+\infty\,.$
Then
$\|\nabla u_{n}\|_{2}^{2}=\delta_{n}\,,$
whereas, recursively integrating by parts,
$\|u_{n}\|_{\ast,p}^{p}$ $=\delta^{\frac{p}{2}}_{n}\ln^{\frac{p}{2}}n\int_{0}^{\frac{R_{n}}{n}}\!\!r\ln(b+r)\,{\rm d}r+\frac{\delta^{\frac{p}{2}}_{n}}{\ln^{\frac{p}{2}}n}\int_{\frac{R_{n}}{n}}^{R_{n}}\!r\ln^{p}\!\left(\frac{R_{n}}{r}\right)\ln(b+r)\,{\rm d}r$
$\leq\frac{\ln^{\frac{p}{2}}n}{2n^{2}}R_{n}^{2}\ln\left(b+\frac{R_{n}}{n}\right)+\frac{1}{\ln^{\frac{p}{2}}n}\int_{\frac{R_{n}}{n}}^{\frac{R_{n}}{{\rm e}}}\!r\ln^{p}\left(\frac{R_{n}}{r}\right)\ln(b+r)\,{\rm d}r$
$\quad+\frac{1}{\ln^{\frac{p}{2}}n}\int_{\frac{R_{n}}{{\rm e}}}^{R_{n}}r\ln(b+r)\,{\rm d}r$
$\leq\frac{\ln^{\frac{p}{2}}n}{n^{2}}R_{n}^{2}+C_{p}\bigg{(}\frac{R_{n}^{2}}{\ln^{\frac{p}{2}}n}\ln\left(b+\frac{R_{n}}{{\rm e}}\right)$
$\quad+\frac{1}{\ln^{\frac{p}{2}}n}\int_{\frac{R_{n}}{n}}^{\frac{R_{n}}{{\rm e}}}\!r\ln^{p-[p]-1}\!\left(\frac{R_{n}}{r}\right)\ln(b+r)\,{\rm d}r\bigg{)}+\frac{R^{2}_{n}}{2\ln^{\frac{p}{2}}n}\ln\left(b+R_{n}\right)$
$\leq\frac{(\ln n)^{\frac{p+2}{2}}}{n^{2}\ln\ln n}+\frac{C_{p}\ln n}{2\ln^{\frac{p}{2}}n\cdot\ln\ln n}\to 0$
since $p>2$. On the other hand,
$\int_{\mathbb{R}^{2}}u_{n}^{p}{\rm e}^{4\pi u_{n}^{2}}\ln(b+|x|)\,{\rm d}x$ $\geq 2\pi\int_{0}^{R_{n}/n}u_{n}^{p}\,{\rm e}^{2\delta_{n}\ln n}r\ln(b+r)\,{\rm d}r$
$\geq\frac{\pi^{1-\frac{p}{2}}}{2^{\frac{p}{2}}}\ln b\,\delta_{n}^{\frac{p}{2}}\,\frac{(\ln n)^{\frac{p+1}{2}}}{(\ln\ln n)^{2}}\to+\infty\quad\text{ as }n\to+\infty\,.$
In view of Theorem 2.5, the next two results are direct consequences.
Proposition 2.6.
The space $X:=D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ is continuously embedded in $L^{q}_{\omega}(\mathbb{R}^{2}):=\{u\in\mathcal{M}(\mathbb{R}^{2})\,|\,\int_{\mathbb{R}^{2}}|u|^{q}\ln(b+|x|)\,{\rm d}x<+\infty\}$ for all $q\geq p$.
Proof.
The proof is a trivial consequence of Theorem 2.5 applied to any normalized function $u/\|u\|\in X$, combined with the elementary inequality
$|t|^{q}\leq C(q,p)|t|^{p}{\rm e}^{t^{2}},\ \ t\in\mathbb{R}\,,$
which holds for any $q\geq p$ and for a proper constant $C(q,p)>1$. ∎
Theorem 2.7.
Under the assumptions of Theorem 2.3, for any $\alpha\leq\frac{4\pi}{\alpha_{0}}$ one has
Finally, we also prove a weighted Cao-type inequality in our space.
Theorem 2.8.
For any $\alpha<4\pi$ and $M>0$
$\sup_{\|\nabla u\|_{2}\leq 1,\ \|u\|_{\ast,p}\leq M}\int_{\mathbb{R}^{2}}|u|^{p}{\rm e}^{\alpha u^{2}}\ln(b+|x|)\,{\rm d}x=C(\alpha,M)<+\infty\,.$
Proof.
The proof follows the same approach as the one for Theorem 2.5, relying on the radial change of variable $T$ and on a quantitive estimate of the Dirichlet norm of the function $u_{\eta}$. As in (2.5)
$\int_{\mathbb{R}^{2}}|\nabla u_{\eta}|^{2}$ $\leq\|\nabla u\|_{2}^{2}+\eta^{\frac{2}{p-2}}\|\nabla\chi\|_{2}^{2}+\eta\|\nabla\chi\|^{2}_{\infty}M^{p}$
$\quad+2\left(\eta^{\frac{1}{p-2}}\|\nabla u\|_{2}\|\nabla\chi\|_{2}+\sqrt{\eta}M^{\frac{p}{2}}\|\nabla u\|_{2}\|\nabla\chi\|_{\infty}\right)$
$\leq 1+\eta^{\frac{2}{p-2}}\|\nabla\chi\|_{2}^{2}+\eta\|\nabla\chi\|^{2}_{\infty}M^{p}+2\eta^{\frac{1}{p-2}}\|\nabla\chi\|_{2}+2\sqrt{\eta}M^{\frac{p}{2}}\|\nabla\chi\|_{\infty}$
$\leq 1+C\eta^{\sigma},\quad\sigma=\min\left\{\frac{1}{p-2},\frac{1}{4}\right\}$
if $\eta<M^{-2p}$. Let us then apply the change of variable $T$, obtaining a new function $v_{\eta}$; we easily obtain
$\|\nabla v_{\eta}\|_{2}^{2}<\left(1+\frac{1}{2\ln(1+\frac{1}{\eta})}\right)\|\nabla u_{\eta}\|_{2}^{2}<1+\frac{C}{|\ln\eta|}\,,$
and
$\|v_{\eta}\|_{p}^{2}\leq\left(1+\frac{C}{|\ln\eta|}\right)\|u_{\eta}\|^{2}_{\ast,p}\leq\left(1+\frac{C}{|\ln\eta|}\right)\|u\|^{2}_{\ast,p}\leq\left(1+\frac{C}{|\ln\eta|}\right)M^{2}.$
Then, for any $\eta$ small enough, depending only on $M$ and $\alpha$, we get
$\int_{B^{c}_{2/\eta}}|u|^{p}{\rm e}^{\alpha u^{2}}\ln(b+|x|)\,{\rm d}x$ $\leq\int_{\mathbb{R}^{2}}|u_{\eta}|^{p}{\rm e}^{\alpha u_{\eta}^{2}}\ln(b+|x|)\,{\rm d}x$
$=\int_{\mathbb{R}^{2}}|v_{\eta}|^{p}{\rm e}^{\alpha v_{\eta}^{2}}\,{\rm d}x=\|\nabla v_{\eta}\|_{2}^{p}\int_{\mathbb{R}^{2}}\left(\frac{|v_{\eta}|}{\|\nabla v_{\eta}\|_{2}}\right)^{p}{\rm e}^{\alpha\|\nabla v_{\eta}\|_{2}^{2}\left(\frac{v_{\eta}}{\|\nabla v_{\eta}\|_{2}}\right)^{2}}\,{\rm d}x$
$\leq C(\alpha,\eta,M)$
if $\alpha\|\nabla v_{\eta}\|_{2}^{2}\leq 4\pi$, that is, if $1+\frac{C}{|\ln\eta|}<\frac{4\pi}{\alpha}$, where $\eta$ is now fixed such that $|\ln\eta|>C\left({\frac{4\pi}{\alpha}-1}\right)^{-1}$ and $\eta<M^{-2p}$. ∎
2.3. Useful theorems and inequalities
@@
Throughout the paper, we will make great use of the following well-known results: the Hardy-Littlewood-Sobolev inequality (see [32, Theorem 4.3]), and the Gagliardo-Nirenberg inequality (see [38, eq. (2.2)]).
@@
Let $N\in\mathbb{N}$ and $u\in L^{p}(\mathbb{R}^{N})$ such that $\nabla u\in L^{r}(\mathbb{R}^{N})$, where $q,r\in[1,+\infty]$. Then, there exists a constant $C:=C(N,p,r,\theta)>0$ such that
$\|u\|_{q}\leq C\|\nabla u\|_{r}^{\theta}\|u\|_{p}^{1-\theta}.$
where $\theta$ satisfies $\frac{1}{p}=\theta\left(\frac{1}{r}-\frac{1}{N}\right)+\frac{1-\theta}{q}$.
We also recall an abstract result from [25, Proposition 3.1] (see also [49, Theorem 2.8] for the version with the Palais-Smale condition), which will be needed to prove the existence of bounded Cerami sequences.
@@
then, for every $\varepsilon\in\left(0,\frac{c-a}{a}\right)$, $\delta>0$, and $\gamma\in\widetilde{\Gamma}$ with $\sup_{u\in M}\Psi(\gamma(u))\leq c+\varepsilon$, there exists $u\in\widetilde{X}$ such that
@@ 2.4. Consequences of the assumptions
To end this section, let us point out some immediate consequences of ($f_{1}$)-($f_{5}$), ($f_{sc}$) and ($f_{c}$), which will be of use in our analysis, together with some comments in this regard:
@@
    while, if $(f_{sc})$ is replaced by $(f_{c})$, the upper bound (2.7) holds for $\alpha>\alpha_{0}$ and, moreover,
    (2.8) $0\leq f(s)\leq C\cdot\begin{cases}s^{p-1}&\mbox{for }s\leq s_{0}\,,\\ s^{r}{\rm e}^{\alpha s^{2}}&\mbox{for }s>s_{0}\,;\end{cases}$
@@
    ($f_{4}$) is related to the well-known de Figueiredo-Miyagaki-Ruf condition [23] and is crucial in order to estimate the mountain pass level and gain compactness, see Lemma 4.5. We note here that such an assumption, which goes back to [24], avoids the prescription of a global lower-bound on $F$ of the kind (2.9) but with $C$ large enough: the latter, indeed, is widely used in the literature but is not of practical verification. A condition similar to ($f_{5}$) appears also e.g. in [14, 7, 17, 42].
@@
    Examples of nonlinearities which satisfy the assumptions of Theorem 1.2 are $F(s)=s^{p}$ and $F(s)=s^{p}{\rm e}^{s^{\alpha}}$ with $p>2$ and $\alpha\in(0,2)$ (concerning ($f_{sc}$)), and $F(s)=s^{p}\chi_{\{s<1\}}(s)+s^{q}{\rm e}^{s^{2}}\chi_{\{s\geq 1\}}(s)$ with $q>-2$ (concerning ($f_{c}$)).